KHO1 Problem 14.35 What is the pressure at a depth of 5000 m in the ocean? your answer with the appropriate units. Part B
Submit Request Answer Part C What is the density of seawater at this pressure? Express your answer with the appropriate units. Submit Provide Feedback O Type here to search
Answer #1

answer a) we have the formula

P=Po+\rhogh

we are adding the gauge pressure as well

Po=1.013*105Pa

\rho=1030kg/m3

sp P=1.013*105+1030*9.8*5000m=5.057*107Pa

so answer is 5.057*107Pa or 5.06*107Pa

b) here we have the formula

V/\DeltaV=-P/B

B= bulk modulus of water

so V/\DeltaV=-5.057*107Pa/2.15*109Pa=-0.025

so the answer is -0.025 or -0.026

c) so here the volume decreases by V-0.025V=0.975

so new density=1030kg/m3/0.975=1056kg/m3

so the answer is 1056kg/m3 or 1057kg/m3.